OpenStudy (anonymous):

Lines CD and DE are tangent to circle A shown below. If C to E is 98°, what is the measure of ∠CDE?

OpenStudy (anonymous):

Select one: a. 56° b. 49° c. 131° d. 82°

OpenStudy (anonymous):

The measure of an angle formed by two tangents is equal to one-half of the difference of the two intercepted arcs. In our problem, the two tangents are intercepted by two arcs, CE and major arc CBE.

OpenStudy (anonymous):

An entire circle has 360 degrees. Since arc CE = 98, major arc CBE = 360 - 98 = 262 degrees.

OpenStudy (anonymous):

Therefore, angle D, the angle formed by the two tangents, is equal to: (1/2) (262-98) = 1/2 of 164 = 82 degrees.

OpenStudy (anonymous):

ok now it makes sense

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Welcome.